The equilibrium constant, Kp, for the following reaction is 55.6 at 698 K: H2(g) + I2(g) 2HI(g) Calculate the equilibrium partial pressures of all species when H2 and I2, each at an intitial partial pressure of 1.65 atm, are introduced into an evacuated vessel at 698 K.
Answer #1

           H2(g) + I2(g) ---------------------- 2aT HI(g)

   1.65 atm       1.65atm                      0

    -x                  -x                            +2x

1.65-x               1.65-x                         +2x

        Kp = P^2HI/PH2xPI2

    55.6 = (2x)^2/(1.65-x)x(1.65-x)

   Square root of 55.6 = 2x/1.65-x

7.46 = 2x/1.65-x

12.309-7.46x = 2x

9.46x = 12.309

x= 12.309/9.46 = 1.30

x= 1.30 atm

At equilibrium

Partial pressure of H2= 1.65-x = 1.65-1.30= 0.35 atm

Partial pressureof I2 = 1.65-x = 1.65-1.30 = 0.35 atm

Partial pressure of HI = 2x = 2x1.30 = 2.60 atm.